10. Find equation of the plane through the point (1,-1,-1) and parallel to the plane 5x – 3y – 2z = 6.please explain thank you.
Answer #1

and pleine 5X-37 -22=6. given by Equation of a plane parallel to 52-34-22=6 is 5x+34 +22f = 6 Also it passes through (1,-1,-1